OpenStudy (anonymous):

Question I got wrong--need someone to explain to me how to find the correct answer and what it is.

OpenStudy (anonymous):

ok what is the first thing to do here ?

OpenStudy (anonymous):

so you should square both sides what do you get ?

OpenStudy (anonymous):

squaring both sides give you 3x -2 <= 4

OpenStudy (anonymous):

since squaring a square root is like removing it and squaring 2 is 2^2 and it is 4

OpenStudy (anonymous):

yes its the same

OpenStudy (anonymous):

≤ is like <=

OpenStudy (anonymous):

so from this inequality what do you get ?

OpenStudy (anonymous):

yes.. and this is what you chose but, there is some restriction here.. the domain of x.. since you have a square root which under it there must be non negative numbers we have another equation 3x - 2 >= 0 from here what do you get ?

OpenStudy (anonymous):

good.. now you have to take the intersection between x<=2 and x>=2/3

OpenStudy (anonymous):

so you get 2/3 <= x <= 2

OpenStudy (anonymous):

:)